1) XYZ Corp. manufactures and sells electric staplers for $8 each. If 5,000 units were sold in December, and management forecasts 2% growth in sales each month, the dollar amount of electric stapler sales budgeted for March should be:

2) XYZ Corp. manufactures and sells electric staplers for $8 each. If 5,000 units were sold in December, and management forecasts 2% growth in sales each month, the number of electric stapler sales budgeted for April should be:
Answer #1

1) Calculate budgeted sales

December January February march
Sales unit 5000 5100 5202 5306.04

Budgeted sales for March = 5306*8 = $42448

2) No of Stapler sales in April = 5306*1.02 = 5412 Units